OpenStudy (anonymous):

Express each ratio as a fraction in simplest form. 18:24

hero (hero):

A ratio such as 18:24 can be written in fraction form: \[\frac{18}{24}\] A fraction can be reduced to simplest form if the number in the numerator and denominator have common factors. The goal is to: 1. Find the greatest common factor between both numbers. 2. Re-write the fraction to include the factors of each number 3. Cancel the GCF 4. The remaining fraction should be relatively prime Since \(6 \times 3 = 18\) \(6 \times 4 = 24\) We can re-write \(\dfrac{18}{24}\) as \(\dfrac{6 \times 3}{6 \times 4}\) Notice that 6 is the GCF Now re-write the fraction as a multiplication of two fractions as follows: \[\frac{6}{6} \times \frac{3}{4}\] Notice that \(\dfrac{6}{6} = 1\) Notice that \(\dfrac{3}{4}\) is relatively prime And \(1 \times \dfrac{3}{4} = \dfrac{3}{4}\)

OpenStudy (anonymous):

ohhhhh ok thank you so much

hero (hero):

By the way if you want to know what relatively prime is: http://www.mathwords.com/r/relatively_prime.htm

hero (hero):

The fraction \(\dfrac{3}{4}\) is in simplest form because the numbers 3 and 4 are relatively prime. In other words, the GCF between both numbers is 1. Which means the fraction cannot be simplified any further.

hero (hero):

I hope that helps.

OpenStudy (anonymous):

yes it does thank you

OpenStudy (anonymous):